OpenStudy (anonymous):

What is the solution set for x2 - 16 = 0?

OpenStudy (jhannybean):

\[\sqrt{x^2-16}=0\]

OpenStudy (anonymous):

(x-4)(x+4)

OpenStudy (jhannybean):

so you can square each side to get rid of the square root \[x^2-16 = 0\] put 16 on the other side, \[x^2=16\] and then take the square root to solve for x \[\sqrt{x^2}=\sqrt{16}\] and you will get two answers :)